# AI-Accelerated Flow Simulation: A Robust Auto-Regressive Framework for Long-Term CFD Forecasting

### v3

@@ Acknowledgments
This work was supported by the Ministry of Science and ICT of Korea grant (No. RS-2024-00355857), and the National Research Council of Science & Technology (NST) grant by the Korea government (MSIT) (No. GTL24031-000). R.V. was supported by ERC Grant No. 2021-CoG-101043998, DEEPCONTROL. The views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of European Union or European Research Council.
@@ Appendix B Evolution of the weights in AW3-based multi-step rollout
In Section 4.4, the weights for each loss term in the AW2-based multi-step rollout are visualized in Fig. 10. This section extends that analysis by presenting results from the AW3-based approach, using the same model configuration as in Fig. 10 but applying AW3 instead of AW2. Fig. 15 illustrates these results, revealing that after 1,000 epochs, the model begins to stabilize the weights of the first and last future steps, with the last time-step weight slightly exceeding that of the first. This suggests that AW3, which initially considers both the first and last time-step losses, remains highly effective by gradually shifting its focus to the last time step, ensuring greater accuracy in long-term rollouts.

### v4

@@ Acknowledgments
This work was supported by the Ministry of Science and ICT of Korea grant (No. RS-2024-00355857, No. 2022-0-00969, and No. 2022-0-00986), the National Research Council of Science & Technology (NST) grant by the Korea government (MSIT) (No. GTL24031-000), and the Ministry of Trade, Industry & Energy (RS-2024-00410810, RS-2025-02317327). R.V. was supported by ERC Grant No. 2021-CoG-101043998, DEEPCONTROL. The views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of European Union or European Research Council.
@@ Appendix B Evolution of the weights in AW3-based multi-step rollout
In Section 4.4, the weights for each loss term in the AW2-based multi-step rollout are visualized in Figure 10. This section extends that analysis by presenting results from the AW3-based approach, using the same model configuration as in Figure 10 but applying AW3 instead of AW2. Figure 16 illustrates these results, revealing that after 1,000 epochs, the model begins to stabilize the weights of the first and last future steps, with the last time-step weight slightly exceeding that of the first. This suggests that AW3, which initially considers both the first and last time-step losses, remains highly effective by gradually shifting its focus to the last time step, ensuring greater accuracy in long-term rollouts.